The supply of wheat is given by the following equation: Qws = -6 + 4Pw - 2Pc - P, where Qws is the quantity o wheat supplied, in millions of bushes; Pw is the price of wheat per bushel; Pc is the price of corn per bushel; and Pf is the price of tractor fuel per gallon. When corn sells for $4 a bushel and fuel sells for $2 a gallon: What is the equation for the supply curve?
Answer #1

Solution:

Given the supply of wheat equation:

Qws = -6 + 4Pw -2c - Pf

Supply curve is simply a relation between the quantity and price of a good. So here we get the supply curve equation for wheat simply by substituting for price of corn per bushel and price of fuel per gallon, in the above equation.

With Pc = $4 and Pf= $2, substituting this in the above equation, we get

Qws = -6 +4Pw -2*4 -2

Qws = - 16 + 4Pw

Or Qws = 4Pw - 16. This is the required equation for the supply curve.